### v1

# Toward Storage-Aware Learning with Compressed Data An Empirical Exploratory Study on JPEG

Abstract
On-device machine learning is often constrained by limited storage, particularly in continuous data collection scenarios. This paper presents an empirical study on storage-aware learning, focusing on the trade-off between data quantity and quality via compression. We demonstrate that naive strategies, such as uniform data dropping or one-size-fits-all compression, are suboptimal. Our findings further reveal that data samples exhibit varying sensitivities to compression, supporting the feasibility of a sample-wise adaptive compression strategy. These insights provide a foundation for developing a new class of storage-aware learning systems. The primary contribution of this work is the systematic characterization of this under-explored challenge, offering valuable insights that advance the understanding of storage-aware learning.
keywords:
On-device AI , Mobile AI , Storage-aware learning
journal: ICT Express
1 Introduction
The rapid advancement of technology over the past few decades has fueled the development of a wide range of applications powered by machine learning. Building on this momentum, the evolution of mobile and embedded devices has ushered in a new era where intelligent systems can operate efficiently under stringent resource constraints [1]. This shift has brought machine intelligence closer to the data source, unlocking a new paradigm that moves beyond mere model execution: on-device model training and personalization [2].
This paradigm addresses the inherent heterogeneity of real-world data from diverse contexts, environments, and hardware, which often undermines the performance of generic models. By continuously learning from individual user data, on-device adaptation mitigates these challenges [3], improving accuracy, robustness, and responsiveness for a more personalized user experience.
However, achieving effective local training on-device introduces several practical challenges. While much of the prior work has focused on overcoming computational limitations such as constrained CPU performance, limited RAM, and energy efficiency [1, 3, 4], a less frequently addressed, yet equally critical, issue is the storage burden [5]. While modern mobile devices are increasingly equipped with larger storage capacities (while still not at par with server-scale platforms), continuous data collection from sensors, cameras, and microphones can quickly consume available space [6]. For instance, recording a stereo audio data at 44.1 kHz and 24-bit precision can require over 22 GB per day ($\sim$660 GB per month). When this is compounded with high-resolution image streams or multi-channel sensor data, the demands on storage escalate dramatically; raising concerns on the storage limitations of mobile and embedded devices.
Under such storage constraints, managing excessive data accumulation typically involves two complementary strategies: (1) controlling the quantity of stored data and (2) improving its quality. Quantity-focused approaches prune samples to reduce dataset size, with early methods relying on static heuristics to discard less informative data [7]. However, the limited adaptability of such heuristics has led to more dynamic techniques, including uncertainty-based and representative sampling [8]. Quality-focused approaches, on the other hand, aim to compress individual samples, often using lossy schemes that allow controlled information loss for substantial storage savings [9]. In practice, these two strategies are interdependent, and addressing them in isolation is rarely sufficient for optimal performance.
Given these two orthogonal approaches, in practical applications, storage constraints are not governed by a single factor but rather by a complex interplay between data quantity and quality. For any given storage budget, a fundamental trade-off exists as one can store a large volume of low-quality (highly compressed) data, or a smaller set of high-quality (less compressed) data. Despite this reality, the majority of existing research has tended to address these dimensions in isolation, focusing on either sample reduction (quantity) or compression techniques (quality) without considering their combined effect. To the best of our knowledge, this work is among the first to systematically characterize two underexplored aspects: first, the impact of storage limitations on the model training, and second, the critical need to balance data quality and quantity to optimize learning outcomes under such constraints.
To effectively explore this quantity vs. quality trade-off, it is crucial to recognize that while the benefits of increasing data quantity are well-established in machine learning, the consequences of altering data quality are far more nuanced and underexplored. Thus, a critical first step towards enabling a strategic balance is to characterize how quality degradation (via compression) fundamentally impacts the learning process, and critically, whether this impact is uniform across all data. Accordingly, this paper addresses the following foundational research questions:
$\bullet$ RQ1 (Training under a storage budget): Given a fixed storage budget, is there a definitive priority between data quality and quantity, or is a strategic balance required to optimize performance?
$\bullet$ RQ2 (Impact of compression): What is the precise impact of data compression and the associated degradation in quality on the behavior of deep learning models?
$\bullet$ RQ3 (Differential Sensitivity): Are all data samples equally sensitive to compression, or do they exhibit varying levels of sensitivity?
As an exploratory study, this work investigates these questions through a series of targeted experiments on a standard computer vision dataset. We note that the goal of this work is to identify a novel problem domain and propose a feasible direction for future research in storage-aware learning. By addressing the aforementioned research questions, this paper offers the following contributions:
  • We highlight the necessity of studying model training under practical, storage-limited scenarios, an often-overlooked constraint in on-device AI.
  • We empirically demonstrate that naive strategies, such as uniform data drop or applying a uniform compression rate, are suboptimal for maximizing performance under a fixed storage budget.
  • We reveal individual data samples exhibit differential sensitivity to compression, which empirically shows the feasibility and potential of developing a sample-wise adaptive compression strategy.
2 Related Work
The performance of machine intelligence consistently improves with increase in training data size, model capacity, and computational resources, a principle known as scaling laws [10]. This paradigm creates a critical dilemma for on-device learning, where the very ingredient that fuels advancement (large-scale data) is also the source of an unsustainable storage burden. Unlike cloud environments with virtually unlimited resources, mobile platforms operate under strict storage constraints. As datasets grow to satisfy data-hungry models, the tension between maintaining high performance and adhering to storage limits becomes increasingly pronounced, especially in environments with continuous data collection, where the rapid accumulation of large volumes of raw data quickly overwhelms limited storage capacity [6].
Consequently, traditional approaches that naively collect and store all available data are no longer viable in these settings. This necessitates a fundamental shift towards more efficient data retention strategies that selectively preserve the most informative content. These strategies primarily involve two levers: (1) selectively discarding samples to manage data quantity and (2) applying compression to manage data quality. Compression reduces storage requirements at the cost of information fidelity [9], while data dropping eliminates samples [11]. This creates a complex trade-off between compression-induced quality degradation, outright data elimination, and final model accuracy.
Therefore, an effective storage-aware training pipeline must be both adaptive and performance-sensitive. Ideally, such a system would dynamically adjust its data retention policy to balance compression rates and sample selection in response to available storage budgets and training objectives. By leveraging techniques such as data valuation or model uncertainty estimates, it becomes possible to allocate storage resources more judiciously. Ultimately, designing training procedures that are responsive to storage limitations is critical for enabling robust and scalable machine learning on next-generation mobile platforms.
3 Empirical Analysis
This section presents an empirical analysis designed to address the foundational research questions outlined in Section 1. Through a series of targeted experiments, we investigate the trade-off between data quality and quantity, the impact of data compression on model behavior, and the differential sensitivity of individual samples to quality degradation. To support our analysis, we conduct experiments using a 3-layer convolutional neural network (CNN) with three convolution layers, trained on the CIFAR-10 dataset [12] for 30 epochs. We deliberately chose this standard dataset and simple architecture to create a controlled environment where the effects of data quality and quantity could be clearly isolated; thus, minimizing confounding factors from more complex models or data distributions.
3.1 Training under a storage budget (RQ1)
Figure 1: Model accuracy as a function of storage usage under various data quantity and quality settings. Each line represents a fixed data quantity (10% to 100%), where points along a line indicate increasing data quality with higher storage usage.
[…]
Figure 1 presents the model’s accuracy on the original (uncompressed, QQ=|D||D|=100%) test set as a function of storage usage, across different data quantities. As expected, model accuracy improves as storage usage increases, reflecting the benefit of higher-quality or larger training datasets. However, a key observation is that under a fixed storage budget, the optimal combination of quantity and quality varies. This observation suggests that maximizing either quantity or quality alone does not always yield the best performance.
Figure 2 further illustrates this point by showing the highest achieved accuracy and the corresponding optimal configuration (i.e., the best quantity–quality pair) for a range of storage constraints. As shown, the ideal trade-off between quantity and quality depends heavily on the available storage, which may be restricted by hardware limitations, system-level policies, or application-specific constraints. These findings emphasize the importance of jointly optimizing both dimensions when designing storage-constrained learning systems.
3.2 Impact of Data Compression (RQ2)
Figure 3: 10 samples from the CIFAR10 dataset with different data quality (i.e., compression rate, $Q$).
Figure 4: Test accuracy as a function of test data quality for models trained with different training data qualities (QQ=5% ($\bullet$), 50% ($\blacksquare$), and 100% ($\blacklozenge$)). The shaded areas R1 and R2 denote the regions where the models trained on 5% and 50% quality data, respectively, achieve optimal performance, while the R3 region is where higher-quality models (QQ=50%, 100%) are superior.
Next, to better understand the impact of applying lossy compression to training data, we trained three CNN models using datasets with varying data qualities (i.e., 5%, 50%, and 100%) and evaluated their performance on test samples compressed at different quality levels. Prior to discussing the results, we first illustrate representative samples to provide intuitive insights. Figure 3 presents the first 10 images from the CIFAR10 test dataset at various data qualities. As expected, image details degrade significantly at higher compression rates (i.e., lower quality); notably, at extremely low quality settings (e.g., $Q=1,5\%$), critical semantic information becomes severely distorted or entirely lost.
Figure 4 illustrates the test accuracy across varying test-sample qualities for models trained with different compression qualities. We note that we use $f_{Q}$ and $Acc_{Q}$ to denote the model trained with quality QQ% data and its accuracy, respectively. Here, we identify three distinct regions, shaded accordingly in the figure: R1 (leftmost region, corresponding to low test quality), where the $f_{5}$ outperforms $f_{100}$ ($Acc_{5}>Acc_{100}$; R2 (middle region, corresponding to mid-range test quality), where $f_{50}$ achieves the highest accuracy among all models ($Acc_{50}>Acc_{100},Acc_{5}$); and R3 (rightmost region, corresponding to high test quality), where $f_{5}$ fails to improve further, while the models trained with higher quality data ($f_{50}$ and $f_{100}$) demonstrate performance improvements and superior accuracy ($Acc_{50},Acc_{100}>Acc_{5}$).
Since $f_{5}$ was unable to capture the rich features present in high-quality data during the training phase, it exhibits poor performance in R3. Nevertheless, it is noteworthy that $f_{5}$ still achieves relatively high accuracy (above $60\%$) in R3, despite the substantial loss of semantic information at very low quality levels (as exemplified in Fig. 3). This performance can be attributed to the fact that low-quality features remain inherently embedded within high-quality data, allowing extremely compressed cases such as $f_{5}$ to effectively exploit these minimal yet informative cues.
Conversely, in R1 and R2, we observed that the model trained with original-quality data ($f_{100}$) performed worse compared to other models. Consistent with observations from R3, while $f_{100}$ can learn low-quality features, it predominantly focuses on those features exclusive to high-quality data. As a result, models trained with lower compression quality, specifically $f_{5}$ and $f_{50}$, demonstrated more robust generalization performance in R1 and R2, respectively.
Based on these observations, we summarize three key findings: (1) Features extracted from low-quality data remain informative and useful (as seen in R3); (2) Relying solely on features exclusive to high-quality data can adversely affect robustness and generalization during inference (evident in R1 and R2); and (3) utilizing intermediately compressed data can offer performance advantages (demonstrated in R1 and R2). Motivated by these empirical results, we hypothesize that it is possible to achieve an optimal balance between storage efficiency and model performance by adaptively selecting appropriate compression qualities.
3.3 Differential Sensitivity to Compression (RQ3)
(a) Loss
[…]
Figure 5: Test loss comparison for original (QQ=100) and compressed (QQ=50) data, sorted by the original loss. (a) Per-sample loss values. (b) Magnitude of the loss difference between QQ=100% and QQ=50%
(a) Loss
[…]
Figure 6: Heatmaps of average loss metrics for test samples grouped into five clusters, evaluated across various compression qualities. (a) The average loss per cluster. (b) The average magnitude of the loss difference per cluster, relative to the original (uncompressed) data.
Rephrasing the first finding in Section 3.2, we observe that certain data samples can be accurately classified without relying on features exclusive to high-quality data (R1), while others cannot (R3). This observation naturally raises the following question: “Are all data samples equally robust or sensitive to compression?” To address this research question, we compared the loss values for each original test sample in the CIFAR10 dataset with the loss values obtained when applying a compression rate of 50%.
Figure 5 (a) and (b) show the actual loss values and the magnitude of the loss differences ($|l_{100}-l_{50}|$) for the original test samples compared to those compressed at 50% quality, where $l_{Q}$ denotes the test loss computed for samples compressed at quality level $Q$. For clarity, the sample indices are sorted based on $l_{100}$ to facilitate intuitive understanding. The results clearly demonstrate that samples well-trained by the model (i.e., those exhibiting relatively low loss values) are less sensitive to compression. In contrast, samples that the model has not learned effectively show significantly higher deviations in loss values compared to their well-trained counterparts.
Extending these observations further, we computed the loss differences between the original test samples and their compressed counterparts across various compression qualities (i.e., 1, 5, 10, 20, …, 90). We then employed the K-means clustering algorithm (with $K=5$ clusters for this experiment) to group these samples based on their compression sensitivity. Figures 6 (a) and (b) present heatmaps illustrating the average loss values and average loss differences across different compression qualities, respectively, along with their corresponding cluster assignments.
As Figure 6 (a) shows, with the exception of Cluster 5, loss values generally tend to increase as the quality of test samples degrades, while each cluster exhibits varying magnitudes and initial degradation thresholds. Specifically, Clusters 1 and 2 demonstrate relatively low loss values for the original high-quality test samples, but exhibit significant loss increases when subjected to the lowest-quality compression. Drawing on observations from Section 3.2, we infer that these clusters represent samples that are well-learned but predominantly reliant on exclusive features only in high-quality data.
In contrast, Cluster 3 exhibits slightly higher initial loss values compared to Clusters 1 and 2 but lower than those of Clusters 4 and 5, coupled with a slower increase in loss as quality decreases, indicating robustness toward compression. Cluster 4, characterized by the highest average loss among Clusters 1 through 4, shows the steepest increase in loss values, highlighting its heightened sensitivity to compression. Lastly, Cluster 5 demonstrates a unique inverse trend, with decreasing loss values as compression quality deteriorates from initially high levels. This behavior suggests that these samples were initially poorly learned, with the model becoming less confused when extraneous, high-quality features are removed through increased compression.
To summarize, the experiments revealed that each data sample exhibits distinct levels of sensitivity and robustness toward compression. Specifically, certain samples minimally influence the model’s training effectiveness when compressed, whereas others significantly affect training performance. These findings justify and strongly motivate an adaptive compression strategy that carefully considers the individual importance of samples during the training process.
3.4 Key Takeaways
In summary, our three-part analysis paints a cohesive picture into understanding the impact of data compression and storage-awareness on AI model performance. Specifically, we first established that an optimal quality-quantity balance exists for a given storage budget (RQ1). We then identified that the impact of this quality degradation is non-uniform, with some models gaining robustness (RQ2) and some samples showing greater resilience to compression (RQ3). Taken together, these findings suggest that the optimal balance discovered in RQ1 is not arbitrary but is a direct consequence of this underlying differential sensitivity. This implies that a truly storage-aware system could leverage prior knowledge of sample sensitivity (from RQ3) to intelligently compress data, thereby achieving near-optimal performance for a given budget without an exhaustive search, as sample sensitivity can be estimated beforehand and used to guide compression decisions.
4 Discussion and Future Works
Towards a Practical Adaptive Policy. Our findings strongly motivate a sample-wise adaptive compression policy, but a key practical challenge remains: how to determine a sample’s compression sensitivity without resorting to exhaustive and computationally expensive evaluations. A promising direction for future work is to explore lightweight, predictive proxies that can estimate this sensitivity. For instance, metrics derived from the model itself, such as prediction uncertainty (e.g., entropy) or gradient-based importance scores, could potentially correlate with a sample’s robustness to compression [11, 8]. A practical system could leverage such a proxy to triage incoming data, dynamically assigning a higher compression rate to samples deemed robust while preserving the fidelity of those identified as sensitive. Developing these predictive proxies is a critical next step to translate our empirical insights into deployable, storage-aware learning systems.
Generalization to Other Data Modalities. A key limitation of our study is its focus on computer vision with a single dataset and compression type (JPEG). While this work focused on computer vision to establish our core principles, the concept of balancing data quality and quantity is modality-agnostic. A compelling avenue for future research is extending this framework to other modalities such as audio, video, and sensor data. Such an extension would necessitate investigating domain-specific compression techniques and their interplay with our proposed storage-aware learning strategy, which would, in turn, significantly broaden the impact and applicability of our findings.
Synergy with Other On-Device Learning Paradigms. Our storage-aware framework is inherently suited for systems with continuous data streams, creating a natural synergy with paradigms like Federated Learning (FL), Continual Learning (CL), and Active Learning (AL) [8, 9, 14]. Our approach can enhance these methods by providing a more sophisticated data retention policy, for example, by enabling efficient knowledge preservation in CL or managing client data under heterogeneous storage in FL. Integrating our quality-quantity balancing strategy into these frameworks presents a significant opportunity to develop more robust and efficient on-device intelligence.
Addressing the Training Overhead. While our primary focus is the storage budget, we acknowledge that retaining compressed data, rather than discarding it, may increase computational overhead during training. We posit this is a manageable trade-off, as on-device training often occurs during device idle periods [15]. Moreover, this overhead could be actively mitigated. Future work could explore hybrid policies that combine data compression with selective dropping or employ data consolidation techniques to form more efficient training batches, thereby optimizing the balance between storage efficiency and computational cost [16].
5 Conclusion
This paper’s primary contribution is the formalization and empirical characterization of storage-constrained on-device learning. Our analysis reveals that the optimal trade-off between data quantity and quality is dynamic and that individual samples exhibit differential sensitivity to compression. These findings invalidate naive, one-size-fits-all data retention policies and establish the feasibility of an adaptive, sample-wise approach. By defining this problem domain, we lay the essential groundwork for a new class of storage-aware learning systems.
Acknowledgements
This work was supported by the Korean MSIT (IITP-2024-2020-0-01461 (ITRC) and IITP-2022-0-00420).